# Early Directional Convergence in Deep Homogeneous Neural Networks for Small Initializations

### v1

@@ A.3 Gronwall’s Inequality
$\left\|\mathcal{H}_{n}(\mathbf{X};\mathbf{s}_{n}(t))-\mathcal{H}_{n}(\mathbf{X};\overline{\mathbf{w}}_{n})\right\|_{2}\leq K^{2}_{\epsilon}\|\mathbf{s}_{n}(t)-\overline{\mathbf{w}}_{n}\|_{2}.$ (78)
Next, for all $t\in[0,\overline{T}_{\delta}]$, we have
$\frac{1}{2}\frac{d\left\|{\mathbf{s}}_{n}-\overline{\mathbf{w}}_{n}\right\|^{2}_{2}}{dt}$ $=(\mathbf{s}_{n}-\overline{\mathbf{w}}_{n})^{\top}\dot{\mathbf{s}}_{n}$
$=\frac{1}{\delta^{L-2}}(\mathbf{s}_{n}-\overline{\mathbf{w}}_{n})^{\top}\mathcal{J}_{n}\left(\mathbf{X};\mathbf{s}_{n}\right)^{\top}(\mathbf{y}-\mathcal{H}_{n}\left(\mathbf{X};\mathbf{s}_{n}\right)-\delta^{L}\mathcal{H}_{z}\left(\mathbf{X};\mathbf{s}_{z}\right))$
$\leq\frac{1}{\delta^{L-2}}(\mathbf{s}_{n}-\overline{\mathbf{w}}_{n})^{\top}\mathcal{J}_{n}\left(\mathbf{X};\mathbf{s}_{n}\right)^{\top}(-\delta^{L}\mathcal{H}_{z}\left(\mathbf{X};\mathbf{s}_{z}\right))$
@@
$\leq\delta^{2}\|\mathbf{s}_{n}-\overline{\mathbf{w}}_{n}\|_{2}\|\mathcal{J}_{n}\left(\mathbf{X};\mathbf{s}_{n}\right)\|_{2}\|\mathcal{H}_{z}\left(\mathbf{X};\mathbf{s}_{z}\right)\|_{2}$
$\leq\delta^{2}\epsilon\alpha\|\mathbf{s}_{n}\|_{2}^{L-1}\beta\|\mathbf{s}_{z}\|_{2}^{L}\leq\delta^{2}\epsilon\alpha\tilde{W}_{\epsilon}^{L-1}\beta\tilde{B}_{\epsilon}^{L},$
where the first inequality follows from (57) and that $\left\|{\mathbf{s}}_{n}(t)-\overline{\mathbf{w}}_{n}\right\|_{2}\leq\epsilon\leq\gamma$, for all $t\in[0,\overline{T}_{\delta}]$. The subsequent set of inequalities follow by Cauchy-Schwartz, and using $\|\mathbf{s}_{z}(t)\|_{2}\leq\tilde{B}_{\epsilon}$ and $\|\mathbf{s}_{n}(t)\|_{2}\leq\tilde{W}_{\epsilon}$, for all $t\in[0,\overline{T}_{\delta}].$ Since $\overline{T}_{\delta}\leq T_{\epsilon}$, integrating the above inequality from 00 to $t\in[0,\overline{T}_{\delta}]$ gives us
$\left\|{\mathbf{s}}_{n}(t)-\overline{\mathbf{w}}_{n}\right\|^{2}_{2}\leq\left\|{\mathbf{s}}_{n}(0)-\overline{\mathbf{w}}_{n}\right\|^{2}_{2}+2\delta^{2}\epsilon\alpha\tilde{W}_{\epsilon}^{L-1}\beta\tilde{B}_{\epsilon}^{L}T_{\epsilon}\leq\delta^{2}(1+2\epsilon\alpha\tilde{W}_{\epsilon}^{L-1}\beta\tilde{B}_{\epsilon}^{L}T_{\epsilon})=C_{\epsilon}\delta^{2}.$ (79)
Next, for all $t\in[0,\overline{T}_{\delta}]$, we have
$\frac{1}{2}\frac{d\|\mathbf{s}_{z}-\mathbf{u}\|_{2}^{2}}{dt}=(\mathbf{s}_{z}-\mathbf{u})^{\top}(\dot{\mathbf{s}}_{z}-\dot{\mathbf{u}})$
$=$ $(\mathbf{s}_{z}-\mathbf{u})^{\top}(\mathcal{J}_{z}(\mathbf{X};\mathbf{s}_{z})^{\top}(\mathbf{y}-\mathcal{H}_{n}(\mathbf{X};\mathbf{s}_{n})-\delta^{L}\mathcal{H}_{z}(\mathbf{X};\mathbf{s}_{z}))-\mathcal{J}_{z}(\mathbf{X};\mathbf{u})^{\top}\overline{\mathbf{y}})$
@@
$=$ $(\mathbf{s}_{z}-\mathbf{u})^{\top}(\mathcal{J}_{z}(\mathbf{X};\mathbf{s}_{z})-\mathcal{J}_{z}(\mathbf{X};\mathbf{u}))^{\top}\overline{\mathbf{y}}+(\mathbf{s}_{z}-\mathbf{u})^{\top}\mathcal{J}_{z}(\mathbf{X};\mathbf{s}_{z})^{\top}\left(\mathcal{H}_{n}(\mathbf{X};\overline{\mathbf{w}}_{n})-\mathcal{H}_{n}(\mathbf{X};\mathbf{s}_{n})-\delta^{L}\mathcal{H}_{z}(\mathbf{X};\mathbf{s}_{z})\right)$
@@
$\leq$ $K^{1}_{\epsilon}\|\overline{\mathbf{y}}\|_{2}\|\mathbf{s}_{z}-\mathbf{u}\|_{2}^{2}+\alpha\|\mathbf{s}_{z}(t)\|_{2}^{L-1}\|\mathbf{s}_{z}-\mathbf{u}\|_{2}\left(\|\mathcal{H}_{n}(\mathbf{X};\overline{\mathbf{w}}_{n})-\mathcal{H}_{n}(\mathbf{X};\mathbf{s}_{n})\|_{2}+\beta\delta^{L}\|\mathbf{s}_{z}\|_{2}^{L}\right)$
$\leq$ $K^{1}_{\epsilon}\|\overline{\mathbf{y}}\|_{2}\|\mathbf{s}_{z}-\mathbf{u}\|_{2}^{2}+\alpha\|\mathbf{s}_{z}(t)\|_{2}^{L-1}\|\mathbf{s}_{z}-\mathbf{u}\|_{2}\left(K_{\epsilon}^{2}C_{\epsilon}\delta^{2}+\beta\tilde{B}_{\epsilon}^{L}\delta^{L}\right)$
$\leq$ $K^{1}_{\epsilon}\|\overline{\mathbf{y}}\|_{2}\|\mathbf{s}_{z}-\mathbf{u}\|_{2}^{2}+\alpha\epsilon\tilde{B}_{\epsilon}^{L-1}(K_{\epsilon}^{2}C_{\epsilon}\delta^{2}+\beta\tilde{B}_{\epsilon}^{L}\delta^{L})$
$\leq$ $K^{1}_{\epsilon}\|\overline{\mathbf{y}}\|_{2}\|\mathbf{s}_{z}-\mathbf{u}\|_{2}^{2}+\alpha\epsilon\tilde{B}_{\epsilon}^{L-1}(K_{\epsilon}^{2}C_{\epsilon}\delta^{2}+\beta\tilde{B}_{\epsilon}^{L}\delta^{2}),$
where in the last equality we used $\overline{\mathbf{y}}=\mathbf{y}-\mathcal{H}_{n}(\mathbf{X};\overline{\mathbf{w}}_{n}).$ In the first inequality we used (77) and the Cauchy-Schwartz inequality. The second inequality follows from (78) and (79), and since $\|\mathbf{s}_{z}(t)\|_{2}\leq\tilde{B}_{\epsilon}$, for all $t\in[0,\overline{T}_{\delta}].$ In last inequality we used $\delta\leq 1.$ Then, integrating the above inequality and using the second Gronwall’s inequality in Lemma A.3, we have
$\|\mathbf{s}_{z}(\overline{T}_{\delta})-\mathbf{u}(\overline{T}_{\delta})\|_{2}^{2}\leq 2\alpha\epsilon\tilde{B}_{\epsilon}^{L-1}(K_{\epsilon}^{2}C_{\epsilon}+\beta\tilde{B}_{\epsilon}^{L})\delta^{2}\overline{T}_{\delta}e^{2\overline{T}_{\delta}K^{1}_{\epsilon}\|\overline{\mathbf{y}}\|_{2}}.$ (80)
For the sake of contradiction assume that $\overline{T}_{\delta}<T_{\epsilon}$. Then, by definition of $\overline{T}_{\delta}$, atleast one of the following equalities must hold
$\|\mathbf{s}_{z}(\overline{T}_{\delta})-\mathbf{u}(\overline{T}_{\delta})\|_{2}^{2}=\epsilon^{2}\mbox{ or }\left\|{\mathbf{s}}_{n}(\overline{T}_{\delta})-\overline{\mathbf{w}}_{n}\right\|^{2}_{2}=\epsilon^{2}$
However, from (79), and since $\delta<\overline{\delta}$, where $\overline{\delta}$ satisfies (65), we have
@@
$\left\|{\mathbf{s}}_{n}(\overline{T}_{\delta})-\overline{\mathbf{w}}_{n}\right\|^{2}_{2}\leq C_{\epsilon}\delta^{2}\leq\epsilon^{2}/2,$
and from (80), we have
$\|\mathbf{s}_{z}(\overline{T}_{\delta})-\mathbf{u}(\overline{T}_{\delta})\|_{2}^{2}$ $\leq 2\alpha\epsilon\tilde{B}_{\epsilon}^{L-1}(K_{\epsilon}^{2}C_{\epsilon}+\beta\tilde{B}_{\epsilon}^{L})\delta^{2}\overline{T}_{\delta}e^{2\overline{T}_{\delta}K^{1}_{\epsilon}\|\overline{\mathbf{y}}\|_{2}}$
$\leq 2\alpha\epsilon\tilde{B}_{\epsilon}^{L-1}(K_{\epsilon}^{2}C_{\epsilon}+\beta\tilde{B}_{\epsilon}^{L})\delta^{2}T_{\epsilon}e^{2T_{\epsilon}K^{1}_{\epsilon}\|\overline{\mathbf{y}}\|_{2}}$
$\leq\epsilon^{2}/2,$
which leads to a contradiction. Thus, $\overline{T}_{\delta}=T_{\epsilon}$. Define $\tilde{C}_{\epsilon}=\sqrt{C_{\epsilon}}$, then from (79), for all $t\in[0,T_{\epsilon}]$, we have
$\left\|{\mathbf{s}}_{n}(t)-\overline{\mathbf{w}}_{n}\right\|_{2}\leq\tilde{C}_{\epsilon}\delta,$
and
$\|\mathbf{s}_{z}(t)-\mathbf{u}(t)\|_{2}\leq\epsilon\implies\|\mathbf{w}_{z}(t/\delta^{L-2})\|_{2}\leq\tilde{B}_{\epsilon}\delta,$ (81)
which proves the first part of the theorem. We next prove the second part of the theorem. Define $\overline{T}_{\epsilon}=T_{\epsilon}/\delta^{L-2}$, then from (81), we have $\|\mathbf{w}_{z}(\overline{T}_{\epsilon})/\delta-\mathbf{u}(T_{\epsilon})\|_{2}\leq\epsilon$. Thus, if $\mathbf{u}(t)$ converges to $\mathbf{0}$, then, from (58) we have $\|\mathbf{u}(T_{\epsilon})\|_{2}\leq\epsilon$, which implies

### v2

@@ C.2 Proof of Theorem 3
$\mathbf{0}=\lambda{\mathbf{a}}_{1}{\mathbf{b}}_{1}^{\top}+\nabla_{{\mathbf{W}}_{1}}\left({\mathbf{y}}^{\top}{\mathcal{H}}({\mathbf{X}};\overline{{\mathbf{W}}}_{1},\cdots,\overline{{\mathbf{W}}}_{L})\right),$ (74)
where $\lambda\neq 0$. Let $q_{1:L-1}\coloneqq q_{1}q_{2}\cdots q_{L-1}$. From Lemma 16, we have
$\nabla_{{\mathbf{W}}_{1}}\left({\mathbf{y}}^{\top}{\mathcal{H}}({\mathbf{x}}_{i};\overline{{\mathbf{W}}}_{1},\cdots,\overline{{\mathbf{W}}}_{L})\right)$ $=\sum_{i=1}^{n}\overline{{\mathbf{W}}}_{2}^{\top}\cdots\overline{{\mathbf{W}}}_{L}^{\top}y_{i}{\mathbf{x}}_{i}^{\top}$
$=q_{1:L-1}\sum_{i=1}^{n}{\mathbf{a}}_{1}{\mathbf{a}}_{2}^{\top}{\mathbf{a}}_{2}{\mathbf{a}}_{3}^{\top}\cdots{\mathbf{a}}_{L-2}{\mathbf{a}}_{L-1}^{\top}{\mathbf{a}}_{L-1}y_{i}{\mathbf{x}}_{i}^{\top}/L^{1/4}$
$=q_{1:L-1}\left(\frac{1}{\sqrt{L}}\right)^{L-2}\frac{1}{L^{1/4}}\sum_{i=1}^{n}{\mathbf{a}}_{1}y_{i}{\mathbf{x}}_{i}^{\top}.$ (75)
From eq. 74 and eq. 75, we have
$\mathbf{0}=\lambda{\mathbf{b}}_{1}^{\top}+q_{1:L-1}\left(\frac{1}{\sqrt{L}}\right)^{L-2}\frac{1}{L^{1/4}}\sum_{i=1}^{n}y_{i}{\mathbf{x}}_{i}^{\top}.$ (76)
Now, ${\mathbf{u}}_{*}$ is a KKT point of
$\max_{{\mathbf{u}}}\left(\sum_{i=1}^{n}y_{i}{\mathbf{x}}_{i}^{\top}\right){\mathbf{u}},\text{ such that }\|{\mathbf{u}}\|_{2}^{2}=1/\sqrt{L},$ (77)
if there exists $\lambda^{*}$ such that
$\sum_{i=1}^{n}y_{i}{\mathbf{x}}_{i}+\lambda^{*}{\mathbf{u}}_{*}=\mathbf{0},\text{ and }\|{\mathbf{u}}_{*}\|_{2}^{2}=1/\sqrt{L}.$ (78)
Hence, from eq. 76, and since $\|{\mathbf{b}}_{1}\|_{2}^{2}=1/\sqrt{L}$, we observe that ${\mathbf{b}}_{1}$ satisfies the KKT conditions in eq. 78.
We now turn towards proving the “if” part.
Note that $(\overline{{\mathbf{W}}}_{1},\cdots,\overline{{\mathbf{W}}}_{L})=({\mathbf{a}}_{1}{\mathbf{b}}_{1}^{\top},q_{1}{\mathbf{a}}_{2}{\mathbf{a}}_{1}^{\top},\cdots,q_{L-2}{\mathbf{a}}_{L-1}{\mathbf{a}}_{L-2}^{\top},q_{L-1}{\mathbf{a}}_{L-1}^{\top}/L^{1/4})$. Now, since ${\mathbf{b}}_{1}$ is a non-zero KKT point of eq. 12, from Lemma 11, there exists $\lambda^{*}\neq 0$ such that
$\sum_{i=1}^{n}y_{i}{\mathbf{x}}_{i}+\lambda^{*}{\mathbf{b}}_{1}=\mathbf{0},$ (79)
We will complete the proof by showing that for $\lambda=q_{1:L-1}\lambda^{*}\left({1}/{\sqrt{L}}\right)^{L-2}/L^{1/4}$,
@@
$\lambda\overline{{\mathbf{W}}}_{l}+\nabla_{{\mathbf{W}}_{l}}\left({\mathbf{y}}^{\top}{\mathcal{H}}({\mathbf{X}};\overline{{\mathbf{W}}}_{1},\cdots,\overline{{\mathbf{W}}}_{L})\right)=\mathbf{0},\forall l\in[L].$ (80)
For $l=1$, we have
@@
$\lambda\overline{{\mathbf{W}}}_{1}+\nabla_{{\mathbf{W}}_{1}}\left({\mathbf{y}}^{\top}{\mathcal{H}}({\mathbf{X}};\overline{{\mathbf{W}}}_{1},\cdots,\overline{{\mathbf{W}}}_{L})\right)$
$=\lambda{\mathbf{a}}_{1}{\mathbf{b}}_{1}^{\top}+\sum_{i=1}^{n}\overline{{\mathbf{W}}}_{2}^{\top}\cdots\overline{{\mathbf{W}}}_{L}^{\top}y_{i}{\mathbf{x}}_{i}^{\top}$
$=\lambda{\mathbf{a}}_{1}{\mathbf{b}}_{1}^{\top}+q_{1:L-1}\sum_{i=1}^{n}{\mathbf{a}}_{1}{\mathbf{a}}_{2}^{\top}{\mathbf{a}}_{2}{\mathbf{a}}_{3}^{\top}\cdots{\mathbf{a}}_{L-2}{\mathbf{a}}_{L-1}^{\top}{\mathbf{a}}_{L-1}y_{i}{\mathbf{x}}_{i}^{\top}/L^{1/4}$
$=\lambda{\mathbf{a}}_{1}{\mathbf{b}}_{1}^{\top}+q_{1:L-1}\left(\frac{1}{\sqrt{L}}\right)^{L-2}\frac{1}{L^{1/4}}\sum_{i=1}^{n}{\mathbf{a}}_{1}y_{i}{\mathbf{x}}_{i}^{\top}$
$=\lambda{\mathbf{a}}_{1}{\mathbf{b}}_{1}^{\top}-q_{1:L-1}\lambda^{*}\left(\frac{1}{\sqrt{L}}\right)^{L-2}\frac{1}{L^{1/4}}{\mathbf{a}}_{1}{\mathbf{b}}_{1}^{\top}=\mathbf{0},$
where the penultimate equality follows from eq. 79. For $12\leq l\leq L-$ 1, we have
@@
$\lambda\overline{{\mathbf{W}}}_{l}+\nabla_{{\mathbf{W}}_{l}}\left({\mathbf{y}}^{\top}{\mathcal{H}}({\mathbf{X}};\overline{{\mathbf{W}}}_{1},\cdots,\overline{{\mathbf{W}}}_{L})\right)$
$=\lambda q_{l-1}{\mathbf{a}}_{l}{\mathbf{a}}_{l-1}^{\top}+\sum_{i=1}^{n}\overline{{\mathbf{W}}}_{l+1}^{\top}\cdots\overline{{\mathbf{W}}}_{L}^{\top}y_{i}{\mathbf{x}}_{i}^{\top}\overline{{\mathbf{W}}}_{1}^{\top}\cdots\overline{{\mathbf{W}}}_{l-1}^{\top}$
$=\lambda q_{l-1}{\mathbf{a}}_{l}{\mathbf{a}}_{l-1}^{\top}+\frac{q_{1:l-2}q_{l:L-1}}{L^{1/4}}\sum_{i=1}^{n}{\mathbf{a}}_{l}{\mathbf{a}}_{l+1}^{\top}{\mathbf{a}}_{l+1}{\mathbf{a}}_{l+2}^{\top}\cdots{\mathbf{a}}_{L-2}{\mathbf{a}}_{L-1}^{\top}{\mathbf{a}}_{L-1}y_{i}{\mathbf{x}}_{i}^{\top}{\mathbf{b}}_{1}{\mathbf{a}}_{1}^{\top}{\mathbf{a}}_{1}{\mathbf{a}}_{2}^{\top}\cdots{\mathbf{a}}_{l-2}{\mathbf{a}}_{l-1}^{\top}$
$=\lambda q_{l-1}{\mathbf{a}}_{l}{\mathbf{a}}_{l-1}^{\top}+\frac{q_{1:l-2}q_{l:L-1}}{L^{1/4}}\left(\frac{1}{\sqrt{L}}\right)^{L-3}{\mathbf{a}}_{l}{\mathbf{a}}_{l-1}^{\top}\sum_{i=1}^{n}y_{i}{\mathbf{x}}_{i}^{\top}{\mathbf{b}}_{1}.$
$=\lambda q_{l-1}{\mathbf{a}}_{l}{\mathbf{a}}_{l-1}^{\top}-\frac{q_{1:l-2}q_{l:L-1}}{L^{1/4}}\lambda^{*}\left(\frac{1}{\sqrt{L}}\right)^{L-2}{\mathbf{a}}_{l}{\mathbf{a}}_{l-1}^{\top}=\mathbf{0},$
where the penultimate equality follows from eq. 79 and since $\|{\mathbf{b}}_{1}\|_{2}^{2}=1/\sqrt{L}.$ Finally, for $l=L$, we have
@@
$\lambda\overline{{\mathbf{W}}}_{L}+\nabla_{{\mathbf{W}}_{L}}\left({\mathbf{y}}^{\top}{\mathcal{H}}({\mathbf{X}};\overline{{\mathbf{W}}}_{1},\cdots,\overline{{\mathbf{W}}}_{L})\right)$
$=\lambda q_{L-1}{\mathbf{a}}_{L-1}^{\top}/L^{1/4}+\sum_{i=1}^{n}y_{i}{\mathbf{x}}_{i}^{\top}\overline{{\mathbf{W}}}_{1}^{\top}\cdots\overline{{\mathbf{W}}}_{L-1}^{\top}$
$=\lambda q_{L-1}{\mathbf{a}}_{L-1}^{\top}/L^{1/4}+q_{1:L-2}\sum_{i=1}^{n}y_{i}{\mathbf{x}}_{i}^{\top}{\mathbf{b}}_{1}{\mathbf{a}}_{1}^{\top}{\mathbf{a}}_{1}{\mathbf{a}}_{2}^{\top}\cdots{\mathbf{a}}_{L-3}{\mathbf{a}}_{L-2}^{\top}{\mathbf{a}}_{L-2}{\mathbf{a}}_{L-1}^{\top}$
$=\lambda q_{L-1}{\mathbf{a}}_{L-1}^{\top}/L^{1/4}+q_{1:L-2}\left(\frac{1}{\sqrt{L}}\right)^{L-2}\sum_{i=1}^{n}y_{i}{\mathbf{x}}_{i}^{\top}{\mathbf{b}}_{1}{\mathbf{a}}_{L-1}^{\top}$
$=\lambda q_{L-1}{\mathbf{a}}_{L-1}^{\top}/L^{1/4}-q_{1:L-2}\lambda^{*}\left(\frac{1}{\sqrt{L}}\right)^{L-1}{\mathbf{a}}_{L-1}^{\top}=\mathbf{0},$
where the penultimate equality follows from eq. 79.
Case 2 $(\alpha\neq 1):$ We begin by proving the “only if” part. Since $\sigma(\cdot)$ is $1-$positively homogeneous, we have
$\sigma(c{\mathbf{p}})=c\sigma({\mathbf{p}}),\text{ and }\sigma(d{\mathbf{q}})={\mathbf{q}}\sigma(d),$ (81)
where $c>0$, ${\mathbf{p}}$ is a vector, and $d\in{\mathbb{R}}$, ${\mathbf{q}}$ is a vector with non-negative entries.
Now, from Lemma 5, for $l\in[L-2]$, we have
$|{\mathbf{a}}_{l}|=|{\mathbf{b}}_{l+1}|,\text{ and }|{\mathbf{a}}_{L-1}|=L^{1/4}|\overline{{\mathbf{w}}}|.$ (82)
Now, since $\{{\mathbf{a}}_{l}\}_{l=1}^{L-1},\{{\mathbf{b}}_{l}\}_{l=2}^{L-1}$ are non-negative, ${\mathbf{a}}_{L-1}=L^{1/4}|\overline{{\mathbf{w}}}|,$ and
${\mathbf{a}}_{l}={\mathbf{b}}_{l+1},\forall l\in[L-2].$
We next show that $q{\mathbf{a}}_{L-1}=L^{1/4}\overline{{\mathbf{w}}}$, for some $q\in\{-1,1\}$. Since $(\overline{{\mathbf{W}}}_{1},\cdots,\overline{{\mathbf{W}}}_{L})$ is a KKT point of eq. 10, from Lemma 16, we have